OpenStudy (anonymous):

-6(4x+3)=6(-4x-3) What is the answer and how do you get it? I am stumped.

OpenStudy (anonymous):

I would start by expanding all that out.

OpenStudy (anonymous):

multiply the number outside the parentheses to each term within the parentheses on both sides of the equation. so do -6 times 4x and -6 times 3 and put the terms back together on the first and do the same thing to the other side of the equation. Then simplify further. You will find that the equation is true and x=x .

OpenStudy (help!!!!):

for this equatin do the distruitive property and you would get -24x-18=-24x-18. The isolate the coefficents and you would get 0

OpenStudy (ryan123345):

First you have to use the distributive property. Then you get -24x-18 = -24x-18 then you do + 24x on each side. Both of them cancel out so you are left with -18 = -18 That is true right? So your answer as always true :)

OpenStudy (anonymous):

are you trying to find x? -6(4x+3)=6(-4x-3) Start with the distributive property, -6*4x=-24x -6*3=-18 so (-24x-18) 6*-4x=-24x 6*-3=-18 so (-24x-18) so -24x-18=-24x-18 add 18 to both sides, -24x-18+18=-24x-18+18 so -24x=-24x divide -24 from both sides, -24/-24 x=0 so now x=0

OpenStudy (anonymous):

\[-6 (4 x+3)=6 (-4 x-3) \]Expand out booth sides.\[-24 x-18=-24 x-18 \] The LHS is equal to the RHS, thus the equation is an identity. x can be any value. " http://www.mathopenref.com/identity.html

OpenStudy (anonymous):

My bad. booth should have been spelled both.

OpenStudy (anonymous):